Crew 187 Sol 9 Summary Report 21JAN2018
Summary Title: Zombie
Author’s name: Cynthia Fuertes Panizo
Mission Status: All nominal
Sol Activity Summary:
Is there someone in Earth?
This day we decided not to do the exercises and start the day late, so at 8:15 we ate breakfast and every Martian started to work in their project, our alarm was delayed an hour. The tiredness begins to be noticed because the naps are more common between meals. At 11:00 our Martian chef, Danton, started to cooked a delicious food: Chicken with cheese, pasta and vegetables. We ate together before the expedition began. At 11:30, Tania, David, Luis and Oscar started the donning time. They took Opportunity and Spirit for this adventure… It’s supposed that they were going to work in El Dorado Canyon, but it was a confusion with the location so they visited a near canyon. However, they took some samples of the halophiles needed. The path that they follow was bent. The canyon that the Martians explorers visited combines a wide variety of landscapes, from rocks up to 5 meters high, which route land for an approximate of 500 meters between ascents and descents and with different types of soil and frozen water remains.
By 15:00, our chef Danton started to cooked a second lunch for our Martian explorers, a delicious soup and scrambled eggs with chili. They arrived at 15:20 to our home. And 25 minutes later we started to eat together… again!
In the afternoon, we got a refill of 125 gallons from the water (trailer). Now we have water for the rest of our lives in Mars, or at least, we hope so.
David is replanting 12 quinua’s plants on the lysimeter for his project. Meanwhile, I get an advanced programing my augmented reality mobile application. And, as always, Tania is documenting everything for her project.
For dinner, Atila is trying to cook a surprise Martian dinner…. Everybody is waiting to taste it!
Cynthia Fuertes Panizo
Commander Crew 187 – MDRS
Look Ahead Plan:
Another EVA is going to take place. The lucky ones are Luis, Atila, David and Danton. They are going to Copernicus Hwy.
Anomalies in work:
Weather: Temperature: -1°C, Fair weather, Winds speed: NNE 8 KPH, Humidity 43%, Barometer 30.22 in
Crew Physical Status: Healthy
Reports to be filed: Sol Summary, GeenHab Report, Operations Report, EVA Request, EVA Report and HSO Report.
Please sent us distillated water.